OpenStudy (anonymous):

What is the axis of symmetry for the graph of y = 3x2 + 6x − 4?

OpenStudy (anonymous):

i didnt just crash

OpenStudy (anonymous):

You can get the axis of symmetry from putting the equation in the vertex form: y = 3(x^2 + 2x) - 4 y = 3(x^2 + 2x + 1) - 4 - 3 y = 3(x + 1)^2 - 7 so the vertex is (-1, -7) and the axis of symmetry is x = -1.

OpenStudy (anonymous):

also what are the domain and range?

OpenStudy (anonymous):

This is a parabola opening upwards, so that vertex is the low-point for "y". The domain is all allowable "x" values which in this case is all numbers. The range is all "y" values and that is y >= -7.

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @Spartan_Of_Ares

OpenStudy (anonymous):

thanks =-]

OpenStudy (anonymous):

where did you get that pic for your profile

OpenStudy (anonymous):

I got it from searching for fractals . I can't remember where though. I like mathematical art.

OpenStudy (anonymous):

ok im going to useit as my descktop background

OpenStudy (anonymous):

go for it!